OpenStudy (anonymous):

Verify the Identity! ( tan^4 x - sec^4 k) = (1 - 2 sec^2 k) Please help me! I'm sure this is a no brainer, but I have no clue how to solve this, I've tried many times and OS is my last resort, please explain how to solve this, anyone! :( please and thank you :O

OpenStudy (anonymous):

try this:\[\Large \tan^2\alpha-1=\sec^2(\alpha) \]

OpenStudy (anonymous):

\[\Large \tan^2x=\sec^2x-1 \] Squaring both sides and that will do.

OpenStudy (anonymous):

Sorry, should be a plus of course. Here again:\[\Large \tan^2x=\sec^2x+1 \]

OpenStudy (anonymous):

\[ a^2-b^2=(a-b)(a+b) \]

OpenStudy (anonymous):

\[ \tan^2x+1 = \sec^2x \]

OpenStudy (anonymous):

yes you're right @wio, I messed this one up.

OpenStudy (anonymous):

@Ikushinkan Still need help?

OpenStudy (anonymous):

is this really an identity with two variables "x" and "k" ???

OpenStudy (anonymous):

@spacelimbus You get the identity wrong; it's: tan^2(x) + 1 = sec^2(x)